### v1

# LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena

Abstract
Recent advancements in large language models (LLMs) have significantly improved performance in natural language processing tasks. However, their ability to generalize to dynamic, unseen tasks, particularly in numerical reasoning, remains a challenge. Existing benchmarks mainly evaluate LLMs on problems with predefined optimal solutions, which may not align with real-world scenarios where clear answers are absent. To bridge this gap, we design the Agent Trading Arena, a virtual numerical game simulating complex economic systems through zero-sum games, where agents invest in stock portfolios. Our experiments reveal that LLMs, including GPT-4o, struggle with algebraic reasoning when dealing with plain-text stock data, often focusing on local details rather than global trends. In contrast, LLMs perform significantly better with geometric reasoning when presented with visual data, such as scatter plots or K-line charts, suggesting that visual representations enhance numerical reasoning. This capability is further improved by incorporating the reflection module, which aids in the analysis and interpretation of complex data. We validate our findings on NASDAQ Stock dataset, where LLMs demonstrate stronger reasoning with visual data compared to text. Our code and data are publicly available at https://github.com/wekjsdvnm/Agent-Trading-Arena.git.
LLM Knows Geometry Better than Algebra:
Numerical Understanding of LLM-Based Agents in A Trading Arena
Tianmi Ma1, Jiawei Du3,4,, Wenxin Huang2, ✉, Wenjie Wang5, Liang Xie6, Xian Zhong1, ✉, and Joey Tianyi Zhou3,4 1 Hubei Key Laboratory of Transportation Internet of Things, Wuhan University of Technology 2 Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University 3 Centre for Frontier AI Research, Agency for Science, Technology and Research, Singapore 4 Institute of High Performance Computing, Agency for Science, Technology and Research, Singapore 5 School of Computing, National University of Singapore  6 School of Science, Wuhan University of Technology
footnotetext: Email:wenxinhuang_wh@163.com, zhongx@whut.edu.cn. denotes the equal contribution. ✉ represents the corresponding author.
1 Introduction
Recent advancements in large language models (LLMs) have demonstrated exceptional proficiency across various domains, achieving state-of-the-art performance in natural language processing (NLP) tasks such as translation (Koshkin et al., 2024), summarization (Yang et al., 2023), and reasoning (Kalyanpur et al., 2024; Qiao et al., 2023). While LLMs excel in language-based tasks, further progress in numerical and geometric reasoning is essential to tackling complex, interdisciplinary challenges, particularly in fields like finance and scientific research. Benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been developed to assess and improve the mathematical problem-solving abilities of LLMs through structured datasets and standardized evaluation protocols. These benchmarks not only advance LLM development but also represent a crucial step in enhancing the ability of artificial intelligence systems to solve mathematical reasoning problems.
Figure 1: Illustration and Performance Comparison of Textual and Visual (K-Line Chart) Numerical Inputs. Top: The left figure presents numerical data in K-line chart format, while the right figure shows textual numerical inputs. Bottom: LLMs with visual inputs (orange) significantly outperform those with textual inputs (blue) in 40-day yield trends, demonstrating the advantage of visual data in financial decision-making.
In response, recent updates to LLMs have focused on improving their mathematical abilities, with models (Ahn et al., 2024; Romera-Paredes et al., 2024) achieving new records on math-focused benchmarks. However, many of these benchmarks (Taylor et al., 2022; Naveed et al., 2023) primarily evaluate performance on problems familiar to the models, closely resembling those encountered during training and often following standard formats, solution strategies, or recurring mathematical patterns. This raises concerns about whether LLMs genuinely possess reasoning abilities for novel numerical data or whether their success is largely due to memorization and pattern recognition. These limitations underscore the need for alternative evaluation paradigms that assess LLMs’ ability to generalize mathematical principles to unseen scenarios.
To address these limitations, we introduce the Agent Trading Arena, a virtual numerical game generating diverse numerical data through agent interactions in a zero-sum stock market environment. In this system, LLM-based agents make trading decisions based on historical stock prices, dynamically determined by agents’ bidding activities. Unlike static benchmarks (Naveed et al., 2023) with predefined strategies, our simulation forces agents to adapt to evolving market conditions. Any temporarily optimal strategy is quickly countered, fostering an environment that rewards adaptability. By mimicking competitive trading scenarios where outcomes depend on real-time agent decisions, we create a controlled yet dynamic environment that tests LLMs’ ability to infer hidden patterns and generalize numerical laws without relying on memorization or prior knowledge.
Our experiments reveal that LLMs struggle with textual numerical data, often focusing on absolute values rather than capturing percentage changes and relations between data points, akin to algebraic reasoning. Additionally, LLMs tend to overemphasize recent data while overlooking earlier information, even when explicitly highlighted in the prompt. These limitations suggest that LLMs lack the ability to abstract numerical information into higher-level representations, a crucial skill for generalizing beyond explicit values.
In contrast, LLMs perform better when presented with numerical data in visualized formats, such as scatter plots, line charts, and bar graphs, which involve geometric reasoning. Our experiments show that LLMs processing visual numerical data consistently outperform those processing textual numerical data. As shown in Figure 1, visual numerical data enables LLMs to achieve significantly higher return rates in the Agent Trading Arena, underscoring the advantages of structured visual representations.
Incorporating the reflection module (Zhou et al., 2022; Shinn et al., 2023) further emphasizes the performance gap between textual and visual data, with the module proving particularly effective for visual representations. By leveraging structured visualizations, the reflection module enhances the model’s ability to reason through complex data relations, leading to more accurate and strategic decision-making.
The Agent Trading Arena provides a valuable framework for advancing LLM generalization and adaptability, with broad applications in fields like finance, healthcare, and scientific research. To further validate our findings, we extend our evaluation to real-world financial data, such as NASDAQ Stock dataset, which includes historical trading data. The consistent results suggest an important contribution: LLMs exhibit stronger reasoning with visual geometric data compared to textual numerical data, demonstrating the value of visual representations in enhancing LLM performance.
2 Related Works
2.1 Mathematical Benchmarks for LLMs
Math word problems (MWPs) have been widely studied, leading to the development of various benchmarks for evaluating models’ mathematical reasoning and problem-solving abilities. Early datasets, such as MAWPS (Koncel-Kedziorski et al., 2016), standardized existing problems to facilitate consistent evaluation. Math23K (Wang et al., 2017) introduced a large-scale collection of Chinese arithmetic problems that require structured equation solving. To increase diversity, benchmarks like ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021) provide richer annotations and a broader range of problem types. More recent benchmarks, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), focus on multi-step reasoning and advanced mathematical concepts, broadening the scope of evaluation. Additionally, MathQA-Python (Austin et al., 2021), a Python variant of MathQA (Amini et al., 2019), emphasizes programmatic reasoning, while MGSM (Shi et al., 2023) extends these benchmarks to multilingual contexts. Despite these advancements, current models (Lu et al., 2023b, a) primarily rely on memory-based answering strategies learned through extensive training, rather than demonstrating true mathematical reasoning.
2.2 LLMs for Enhanced Mathematical Reasoning
Building on these benchmarks, LLMs have advanced mathematical problem-solving by incorporating specialized datasets into their training. Models such as Galactica (Taylor et al., 2022), PaLM-2 (Anil et al., 2023), Minerva (Lewkowycz et al., 2022), and LLaMA-2 (Touvron et al., 2023) leverage extensive datasets during pre-training, improving their mathematical reasoning abilities and understanding of complex concepts. Fine-tuned models like MetaMath (Yu et al., 2024), MAmmoTH (Yue et al., 2024), and WizardMath (Luo et al., 2023) are specifically tailored for mathematical tasks. These models undergo domain-specific fine-tuning with carefully curated datasets, enabling them to tackle advanced reasoning challenges and significantly enhance performance on specialized benchmarks.
However, existing methods (Naveed et al., 2023; Yang et al., 2024a) often rely on large training datasets, raising concerns about the true reasoning capabilities of LLMs. While these methods primarily assess performance on established problem types, the heavy reliance on extensive training data suggests that models may achieve high performance through memorization and pattern recognition, rather than genuine reasoning. Consequently, alternative evaluation paradigms are needed to more accurately assess LLMs’ ability to generalize mathematical principles to novel scenarios.
Figure 2: Stock Trading Workflow in Agent Trading Arena. Top: Workflow of a trading day, including preparation, trading, and post-trading reflection. Agents discuss insights in the chat pool, analyze market trends, execute trades, and refine strategies based on performance. Bottom: Example of agents’ interactions in the chat pool and dynamic strategy updates.
3 Proposed Method
To mitigate the influence of human prior knowledge and memory, we designed a closed-loop economic system (Guo et al., 2024) called the Agent Trading Arena, a zero-sum game simulating complex, quantitative real-world scenarios. The simulation workflow is illustrated in Figure 2 and further detailed in Appendix A. In the Agent Trading Arena, agents can invest in assets, earn dividends from holding assets, and pay daily expenses using virtual currency. The agent with the highest total return wins the game.
3.1 Agent Trading Arena
Structure of Agent Trading Arena.
To eliminate external knowledge biases, asset prices are determined by a bid-ask system, reflecting the prices at which buyers and sellers are willing to transact. The system evolves solely based on agents’ actions and interactions, without external influences. This design ensures that the outcomes of agents’ actions are not immediately apparent but unfold gradually, influenced by other agents’ decisions.
To encourage active participation, a dividend mechanism is introduced. There are two primary sources of income in this system: capital gains from asset price differentials and dividends from holding assets. Dividends for each asset are distributed according to a predefined ratio, serving as an implicit anchor for asset prices. Agents holding more low-cost assets receive higher dividends. To prevent passive asset holding until the end of the game, agents must pay a daily capital cost proportional to their total wealth. These expenses are offset by asset dividends, and only agents with sufficient low-cost assets can cover costs. Under the pressure of significant daily expenses, agents must act swiftly and strategically, triggering frequent trades and price fluctuations to stimulate market activity. This dynamic mechanism ensures fairness in the zero-sum game while preventing agents from relying on fixed strategies to find optimal solutions.
Agents Learn and Compete in Arena.
The zero-sum game structure is crucial to eliminating the possibility of a universally optimal strategy. In fixed scenarios with a static optimal solution, agents could rely on predefined rules or memory-based approaches, bypassing adaptive decision-making. The zero-sum game ensures that there is no universally correct solution, with outcomes evolving dynamically based on agent interactions and competition. This design forces agents to continually adapt, learn from feedback, and develop context-dependent strategies, promoting deeper environmental exploration and preventing reliance on static or memory-driven solutions.
In the Agent Trading Arena, agents are unaware of implicit rules, except for the objective to maximize their virtual wealth throughout the simulation. To win this zero-sum game, agents must effectively learn from experience, decipher hidden game rules, and develop strategies to counter competitors. This requires the ability to comprehend numerical feedback, formulate enduring strategies, and make informed decisions. Unlike other mathematical reasoning problems, the results of their actions unfold gradually and dynamically. Moreover, agents are easily misled by erroneous information from competitors, hindering their ability to discern strategic cues from competitors’ textual data. Importantly, agents remain unaware of these implicit rules, so applying real-world knowledge does not benefit their performance. Therefore, agents must rely on experiential learning to decipher the hidden game rules and ultimately achieve victory.
3.2 Types of Numerical Data Input
Limitations of Textual Numerical Data.
In the Agent Trading Arena, the generated stock data is stored in numerical format. When used directly as input to an LLM, the models often struggle to interpret numerical data accurately or make sound decisions. To mitigate this, we convert the data into textual formats (Hu et al., 2024; Zhang et al., 2024), enhancing semantic features and clarifying output requirements to improve the models’ understanding. During interactions, the LLMs process stock prices, trading volumes, and market indices presented as textual numerical data.
Figure 3: Textual and Visual Representations of Corresponding Inputs and Outputs. The left images display the agent’s Buy and Sell trading records, daily trade prices, and K-line charts for three stocks. The output from visual inputs (bottom right) captures overall stock trends and long-term behavior, while the output from textual inputs (top right) focuses on specific current prices.
However, this textual approach reveals significant limitations. While the data is presented clearly, LLMs tend to focus excessively on specific values rather than identifying long-term trends or global patterns. They also struggle with understanding correlative relations and percentage changes, limiting their ability to assess differences and identify connections between data points. When analyzing time-series data with complex patterns, LLMs often fixate on individual data points, overlooking overarching relations. This issue is evident in the analysis output in the top-right corner of Figure 3, where LLMs’ focus on individual values impedes their ability to generalize, reducing their capacity to extract meaningful global insights.
Additionally, LLMs often overemphasize recent data while undervaluing historical information, even when prompted to consider its importance. This prevents them from effectively integrating past data and recognizing long-term patterns, complicating their understanding of numerical relations and trends. These challenges highlight the need for improved mechanisms to process numerical relations, identify global trends, and derive deeper insights from textual numerical data.
Potential of Visual Numerical Data.
Since textual numerical data often leads LLMs to focus on local details while neglecting broader relations, we investigated whether visual representations, such as scatter plots, line charts, and bar charts, could help LLMs better understand overall trends, similar to human reasoning. Thus, we transition from textual numerical data inputs to visualized formats  (Yang et al., 2025). As demonstrated in the bottom-right corner of Figure 3, visual representations enable LLMs to more effectively grasp global trends, patterns, and relations that are often difficult to discern from textual numerical data alone.
These findings highlight the advantages of structured, visual numerical data, indicating that this format allows LLMs to more intuitively and comprehensively understand complex data, better capturing overall fluctuations, whereas text tends to focus on local details. By combining visualization and textual representations, LLMs not only overcome the challenges of relations in time-series data but also demonstrate better performance in identifying long-term trends and global patterns, while still attending to local details.
3.3 Reflection Module
We propose a strategy distillation method, illustrated in Figure 4, that delivers real-time feedback to LLMs by analyzing both descriptive textual and visual numerical data. This enables the generation of new strategies and optimization of action plans. The approach allows agents to evaluate their results, refine strategies, and adapt continuously based on feedback. The process begins with assessing the day’s trajectory memory and associated strategies using an evaluation function. The strategic generation process leverages contrastive analysis of peak and nadir performers from the evaluation phase, creating bidirectional learning signals that inform subsequent iterations. This iterative cycle ensures continuous strategy evolution, fostering sustained improvement in decision-making.
Figure 4: Design of the Reflection Module. The process evaluates daily trajectory memory and strategies (top right), then generates new strategies (center) based on evaluation, environmental feedback (bottom right), and feedback from the 5 top- and bottom-performing strategies. Stock visualization (bottom left) enhances reflection, driving continuous improvement.
The reflection module plays a crucial role in refining strategies by offering real-time feedback. It analyzes both descriptive textual and visual numerical data to generate new strategies and optimize action plans. Within the Agent Trading Arena, the reflection module is triggered regularly to consolidate daily trading records and evaluate the effectiveness of strategies, refining both successful and unsuccessful experiences to guide future decisions. Ineffective strategies are stored in a strategy library for future reference, allowing agents to review and learn from past experiences. Further details can be found in Appendix A.
ID Name Year Cash Occupation
1 Amy 1 100,000 AI Researcher
2 Bruce 2 100,000 Lawyer
3 Charles 1 100,000 Doctor
4 David 3 100,000 Engineer
5 Ella 2 100,000 Teacher
6 Frank 5 100,000 Entrepreneur
7 Grace 4 100,000 Accountant
8 Hank 2 100,000 Architect
9 Ivy 3 100,000 Marketing Manager
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 1: Agent Details. “Year” is the investment duration, and “Cash” is the initial capital.
ID Name DPS Historical Close Prices Quantity
1 A 22 454.17, $\dots$, 445.60 1,200
2 B 23 354.17, $\dots$, 465.80 1,000
3 C 25 500.47, $\dots$, 440.60 1,600
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 2: Stock Details. “DPS” is the Dividend Per Share, and “Quantity” is the initial share count.
4 Experimental Results
4.1 Experimental Setup
Datasets.
To evaluate the ability of LLMs to analyze and process data, we developed the Agent Trading Arena, a controlled environment that isolates external factors. The system’s workflow is illustrated in Figure 2. Within this environment, agents discuss stock market trends, analyze stock data, and engage in trading activities. Each agent can execute multiple trades per day and reflect on all trades at the end of each trading session.The Agent Trading Arena allows for adjustable numbers of agents and stocks; in our experiment, we deployed at least nine agents and three stocks. All agents were provided with the same initial capital to ensure identical starting conditions. Detailed information about each agent is presented in Table 1, and stock-specific data is available in Table 2. To further validate our findings, we selected a subset of NASDAQ Stock dataset for portfolio investment. For more details, refer to Appendix B.
Evaluation Metrics.
Each agent was assigned varying capital based on their roles, and performance was evaluated using the following metrics:
1) Total Return (TR): Measures the overall performance of the strategy, calculated as: $\mathrm{TR}={(C_{1}-C_{0})}/{C_{0}}$, where $C_{0}$ is the initial asset value and $C_{1}$ is the final asset value.
2) Win Rate (WR): Represents the proportion of winning trades, calculated as: $\mathrm{WR}={N_{w}}/{N_{t}}$, where $N_{w}$ is the number of winning trades and $N_{t}$ is the total number of trades.
3) Sharpe Ratio (SR): Evaluates investment returns relative to risk, calculated as: $\mathrm{SR}={(R_{p}-R_{f})}/{\sigma_{p}}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is the risk-free return, set to 0 as defined in SocioDojo (Cheng and Chin, 2024).
4) Mean Daily Return (Mean): Indicates the average return per day during the trading period.
5) Standard Deviation of Daily Returns (Std): Reflects the volatility of daily returns, indicating the risk associated with the strategy.
4.2 Comparative Experiments
We conducted experiments to assess the real-time data analysis and reasoning capabilities of LLM-based agents, focusing on how textual and visual representations influence decision-making. First, we explored the impact of textual and visual representations in dynamic environments. Next, we incorporated the reflection module to enhance agents’ reasoning and data interpretation, examining how reflective reasoning influences decision-making. Finally, we validated the model’s effectiveness through stock investment simulations on NASDAQ Stock dataset, assessing agents’ adaptability and decision-making in real-world scenarios.
LLMs Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
w/o Reflection LLaMa-3 (Dubey et al., 2024) 4.0934 0.6857 2.2700 62.5001 0.1963
DeepSeek (Liu et al., 2024) 27.3078 2.2478 2.7553 92.2078 1.5303
Qwen-2.5 (Yang et al., 2024b) 30.3740 2.6795 1.4588 93.7500 1.7025
Gemini-1.5 (Reid et al., 2024) 14.3193 1.2210 2.7776 83.8384 1.7293
19.0389 1.6038 1.1654 90.9091 1.3761
23.9649 1.9809 1.3342 100.0000 1.4847
GPT-4o (Hurst et al., 2024) 13.0369 1.1386 1.9285 54.5455 0.5904
17.0661 1.4539 1.5040 72.7272 0.9668
26.1806 2.1574 2.0578 90.9091 1.0484
w/ Reflection LLaMa-3 (Dubey et al., 2024) 10.2458 1.6669 2.4034 66.6667 0.6936
DeepSeek (Liu et al., 2024) 30.6238 2.4696 1.5121 100.0000 1.6332
Qwen-2.5 (Yang et al., 2024b) 38.9113 3.3244 1.0726 100.0000 3.0994
Gemini-1.5 (Reid et al., 2024) 29.4511 2.5185 3.0116 93.5064 2.2100
37.0054 2.9111 1.2349 100.0000 2.3574
41.3264 3.1946 0.1588 100.0000 20.1128
GPT-4o (Hurst et al., 2024) 33.6508 2.6713 2.1711 98.7013 2.1417
35.7622 2.8206 0.6782 100.0000 4.1590
47.6851 3.6095 0.5327 100.0000 6.7765
Table 3: Performance Comparison Using Textual and Visual Data w/o and w/ Reflection. Evaluation of agent portfolios show that visual approaches outperformed textual ones, with the best results achieved by combining both. LLMs demonstrated a significant advantage in TR and SR, with visual inputs yielding greater improvements than textual inputs. The best and second-best results are highlighted in bold and underlined.
Trials with Textual or Visual Input.
To enhance LLMs’ understanding of complex data in the Agent Trading Arena, we transitioned from textual numerical inputs to visualized formats, including scatter plots, line charts, and bar charts. Three types of visualizations were used: daily K-line charts, transaction histories, and quantities traded by each agent. The experimental setup for the visualization group is shown in Figure 3 and detailed in Appendix C. In the Arena, LLMs without image input capabilities received only text input and did not perform reflection. For image-enabled LLMs, the first agent received only visual input, the second received both textual and visual input, and the others received only textual input. None of the agents had reflection capabilities. For details on the selected LLMs, please refer to Appendix E.
We conducted experiments across different LLMs and the results for Gemini-1.5 (Reid et al., 2024) and GPT-4o (Hurst et al., 2024) in Table 3 show that agents with visual numerical input outperformed those with textual input alone. Agents receiving both textual and visual input achieved the best performance. This combination enables agents to focus on local details while also understanding overall trends, leading to optimal performance. This suggests that LLMs exhibit better reasoning skills with visual geometric data rather than textual numerical data, highlighting a clear distinction in their strengths in geometry and algebra.
Strategy TR $\uparrow$ SR $\uparrow$
MACD (Chong and Ng, 2008) 7.18 0.173
StockFormer (Gao et al., 2023) 9.05 0.073
TimesNet (Wu et al., 2023) 11.74 0.203
GPT-4o + Textual 8.69 0.167
GPT-4o + Visual 9.91 0.195
GPT-4o + Textual + Visual (Ours) 12.23 0.291
Table 4: Reproduced Backtesting Performance on NASDAQ Stock Dataset. The agent with textual and visual input outperformed the NASDAQ-100 by 53.97% in SR during the same period.
LLMs T V TR $\uparrow$ SR $\uparrow$
LLaMa-3 100.00 100.00
DeepSeek + 35.03 + 45.63
Gemini-1.5 + 10.99 + 14.93
+ 11.49 + 29.47
GPT-4o + 10.39 + 28.29
+ 17.18 + 40.77
Table 5: Performance Comparison in Trading Decisions w/o Reflection, Competing Pairwise with LLaMa-3. DeepSeek may possess unique strengths or optimizations that allow it to better adapt to the task’s complexities when competing with other models.
LLMs TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
LLaMa-3 (Dubey et al., 2024) 100.0000 100.0000 100.0000 100.0000 100.0000
Gemini-1.5 (Reid et al., 2024) + 3.8631 + 5.9249 + 12.0426 - 4.0010 + 2.2920
DeepSeek (Liu et al., 2024) + 3.1317 + 3.1309 + 1.7169 + 9.8039 + 1.7945
Qwen-2.5 (Yang et al., 2024b) + 7.4578 + 7.3907 - 2.4801 + 50.0000 + 8.0524
GPT-4o (Hurst et al., 2024) + 17.7636 + 16.9041 + 2.4136 + 37.9310 + 14.2775
Table 6: Performance Comparison in Trading Decisions w/ Reflection, Competing Pairwise with LLaMa-3 Using Textual Data. The results show that in competition with reflection using textual numerical data, GPT-4o and Qwen-2.5 outperformed other models.
Trials with Reflection and Textual or Visual Input.
Building upon the previous experiments, we incorporated the reflection module to investigate its impact on the experimental outcomes. By introducing this module, we aimed to determine whether reflective processing could improve strategies and adaptability in complex financial scenarios.
Our experiments with diverse LLMs demonstrate that reflective agents consistently outperformed non-reflective agents in stock trading. As shown in Table 3, text-only LLMs with reflection modules achieved enhanced returns, with GPT-4o (Hurst et al., 2024) reaching 33.65% total return using textual input alone, which increased to 47.70% with visual data. Combining both input types resulted in a 41.74% higher return compared to using textual input alone. Compared to agents without the reflection module, including the reflection module significantly enhanced each agent’s stock investment performance. LLMs with the reflection module showed the greatest improvement in agents with visual inputs, compared to those with textual inputs, in both the Total Return and Sharpe Ratio. This indicates that incorporating reflection further widens the gap between LLMs’ geometric reasoning with visual data and algebraic reasoning with textual numerical data, highlighting their strengths in both areas.
Simulation with NASDAQ Stock dataset.
To further validate our findings, we conducted a two-month investment simulation on NASDAQ Stock dataset, starting with an initial capital of 100,000 units for portfolio investment. NASDAQ Stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. Detailed information about the models used is provided in Appendix B.
The results in Table 4 show that, despite StockFormer (Gao et al., 2023) and TimesNet (Wu et al., 2023) benefiting from longer training periods and larger datasets, they underperformed compared to our model. Notably, our model achieved superior results without additional training, relying solely on historical stock data for decision-making. The Sharpe ratios for Apple, NASDAQ-100, and S&P 500 during the same period were 0.097, 0.189, and 0.205, respectively. The agent with textual and visual input outperformed the NASDAQ-100 and S&P 500 in Sharpe Ratio by 53.97% and 41.95%, respectively. Moreover, only visual input outperformed textual input, further highlighting LLMs’ stronger geometric reasoning abilities.
4.3 Ablation Experiments
Impact of Modality on LLM Competitiveness.
We employed a relative evaluation method for this experiment. The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline. This setup aimed to explore the impact of different agents and modalities on LLM performance. The results are shown in Table 5. The findings indicate that DeepSeek (Liu et al., 2024) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
Competition Among Different LLMs with Reflection.
To investigate the role of reflection in competition dynamics using textual numerical data between LLMs, we conducted an ablation study in the Agent Trading Arena. Using a relative evaluation approach, the first agent employed various LLMs, while eight LLaMa-3-based agents served as the baseline for a fair comparison. This setup effectively isolated the influence of each LLM on performance. As shown in Table 6, in competition with reflection, GPT-4o (Hurst et al., 2024) and Qwen-2.5 (Yang et al., 2024b) outperformed other models, consistent with the findings in Table 3.
5 Conclusion
In this paper, we introduced the Agent Trading Arena, a zero-sum game designed to simulate complex economic systems and evaluate LLMs on numerical reasoning tasks. Our results show that while LLMs struggle with plain-text numerical data (algebraic reasoning), their performance significantly improves when presented with visual data (geometric reasoning). This highlights the advantage of visual representations in supporting numerical reasoning, particularly in complex scenarios. The integration of a reflection module further enhances model performance, allowing LLMs to analyze better and interpret data. We validated these findings on NASDAQ Stock dataset, demonstrating that LLMs excel in visual geometric reasoning tasks, suggesting that LLMs may perform better with visual numerical data than with textual numerical data. Overall, our work offers insights into the strengths and limitations of LLMs in dynamic numerical reasoning tasks, particularly in the context of geometry vs. algebra, and sets the foundation for future research on improving their performance in real-world, interdisciplinary challenges.
Limitations
This study evaluates LLMs within a virtual stock trading environment, focusing on their performance in visual geometric reasoning. While this controlled setting limits the generalizability to other domains, it provides valuable insights into LLMs’ capabilities. The reliance on high-quality visualizations, reflection modules, and substantial computational resources may restrict applicability in resource-constrained environments. Future research can address these limitations by broadening the scope to include diverse reasoning tasks, optimizing computational requirements, and exploring alternative modalities for more robust and generalizable assessments.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University under Grant 2024BDIAA01.
Appendix A Agent Trading Arena
A.1 Agent Trading Arena Details
As illustrated in Figure 5, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection. In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
Figure 5: Agent Workflow Components: Share Trading involves stock analysis, decision validation, and trade execution; Environment and Memory manage memory and process trade orders; and Reflection focuses on strategy assessment and refinement based on feedback.
A.1.1 Action Decision-Making
Action generation follows the LLM framework. The agent responsible for generating actions receives corresponding prompts via SQLite. Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks. The action generation process is outlined in Equation 1, with input prompts shown in Figure 6 and the corresponding outputs displayed in the adjacent figure.
$\begin{cases}A_{\mathrm{date}+1}^{t}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{if~{}}t\text{~{}is Iters},\\ A_{\mathrm{date}}^{t+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{otherwise},\end{cases}$ (1)
where $\mathrm{Ins}$ represents the environment introduction, $Z_{\mathrm{date}}^{t}$ denotes the memory of the stock transaction on day $\mathrm{date}$ retrieved from the database, and $S_{\mathrm{date}}$ is the strategy for day $\mathrm{date}$ generated via reflection.
Figure 6: Inputs and Outputs in Action Decision-Making.
A.1.2 Environmental Interaction
To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment. The function $\phi$ facilitates environmental interactions, as shown in Algorithm 1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment. Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares. Stock prices are independent of external factors and are influenced solely by the sandbox’s internal dynamics. The stock price is updated with each transaction according to the following formula:
$\mathrm{Price}_{\mathrm{curr}}$ $=\delta\left(Q,F,\mathrm{Price}_{\mathrm{curr}},\mathrm{Price}_{\mathrm{deal}}\right)$ (2)
$=\frac{\mathrm{Price}_{\mathrm{deal}}\cdot Q\cdot F+\mathrm{Price}_{\mathrm{curr}}\cdot Q_{\mathrm{total}}}{Q\cdot F+Q_{\mathrm{total}}},$
where $Q$ is the quantity of stock traded, $F$ is the fluctuation constant, $\mathrm{Price}_{\mathrm{curr}}$ is the current stock price, $Q_{\mathrm{total}}$ is the total number of shares available, and $\mathrm{Price}_{\mathrm{deal}}$ is the price at which the trade occurs.
Input: OPS: Function to retrieve agents’ actions, date: The current trading date, Z: Memory used for interaction with the environment
Output: Z
for $P\in\text{Persons}$ do
       $A\leftarrow\text{OPS}(t,P)$;
       $O,N,Q,\text{Price}_{\text{deal}}\leftarrow\text{Extract}(A)$;
       $\text{Price}_{\text{curr}},Q_{\text{total}}\leftarrow\text{Stocks}(N)$;
       $\text{Price}_{\text{curr}}\leftarrow\delta(Q,F,\text{Price}_{\text{curr}},\text{Price}_{\text{deal}})$;
      if $O=$ “buy” then
             $\text{Cash}\leftarrow\text{Price}_{\text{curr}}\cdot Q$;
             if $<P.\text{Cash}$ Cash then
                   $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
      if $O=$ “sell” then
             $\text{Hold}\leftarrow P(N)$;
[…]
                   $Q_{N}\leftarrow\text{Hold}[\text{``Q''}]-Q$;
                   if $Q_{N}>0$ then
                         $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
[…]
return Z;
Algorithm 1 Environmental Interaction
The function $\phi$ executes trading orders and updates the stock price in real-time based on the agent’s actions. To prevent excessive volatility and mitigate risk during trading cycles, a daily price fluctuation cap is enforced. Before executing any transaction, each agent evaluates its available funds and refrains from proceeding if insufficient capital is available.
A.1.3 Memory
The superior performance of LLM-based agents arises from the extensive internal knowledge acquired during pre-training. The large number of parameters in LLMs enables the retrieval of diverse information and supports logical and inferential reasoning. To further enhance knowledge retrieval across various tasks, we incorporate a memory module that empowers LLM-based agents with self-improvement capabilities. This memory module facilitates strategy reflection through time-series feedback. Unlike qualitative tasks, quantitative feedback evolves incrementally with subtle differences, presenting a challenge for the generalization of existing LLM-based agents.
To minimize the influence of pre-existing knowledge, we assign specific roles to each agent. The memory module records sensory inputs and accumulates valuable experiences based on immediate feedback following actions. These experiences are stored in a database for future reference. The interaction history between an agent’s behavior and the environment constitutes short-term memory, enabling the agent to retain recent events. The historical trajectory is defined as:
$M_{\mathrm{date}}^{t}=\zeta\left(I_{\mathrm{date}}^{t},\mathrm{Out}_{\mathrm{date}}^{t}\right),$ (3)
where $\zeta$ processes key information, $I_{\mathrm{date}}^{t}$ represents the input prompt, and $\mathrm{Out}_{\mathrm{date}}^{t}$ denotes the resulting output.
The day’s trajectory constitutes short-term memory, which is expressed as:
$Z_{\mathrm{date}}^{t}=\left(M_{\mathrm{date}}^{0},M_{\mathrm{date}}^{1},\dots,M_{\mathrm{date}}^{t}\right).$ (4)
The process for updating short-term memory is given by:
$Z_{\mathrm{date}}^{t+1}=M_{\mathrm{date}}^{t}\cup Z_{\mathrm{date}}^{t},\quad t\in\{0,1,\dots,T\},$ (5)
where $Z_{\mathrm{date}}^{t}$ represents the short-term memory, and $T$ denotes the maximum number of iterations.
The reflection model serves as long-term memory, enabling self-reflection and the consolidation of knowledge.
Figure 7: Visualization of Stock Inputs and Corresponding Trading Strategy Outputs.
A.1.4 Reflection
We propose a strategy distillation method that transforms quantitative results into descriptive text, which is then used as prompts for LLMs. This approach aids in the analysis of results and the generation of actionable, qualitative summaries, enabling LLMs to derive new strategies. These strategies are implemented, monitored, and evaluated over time, while underperforming strategies are archived for future review.
Initially, we evaluate the day’s trajectory memory and associated strategies. The evaluation function is defined as:
$E_{\mathrm{date}}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},S_{\mathrm{date}}\right),$ (6)
where $\Psi$ represents the evaluation function based on LLMs, $\mathrm{Ins}$ contains the agent role descriptions and output requirements, and $Z_{\mathrm{date}}^{T}$ denotes the memory for the given day.
Based on the evaluation results and previous strategies, we generate the latest strategy. Past strategies are stored in a library and scored. For the new strategy, we select the top five best-performing and the bottom five worst-performing strategies to provide both positive and negative feedback. The strategy update formula is:
$S_{\mathrm{date}+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{T},E_{\mathrm{date}},S_{\mathrm{date}}\right),$ (7)
where $E_{\mathrm{date}}$ represents the evaluation results for the day, and $S_{\mathrm{date}}$ is the strategy for that day.
Long-term memory is generated through reflection, as represented by:
$Z^{\prime}_{\mathrm{date}}=\left(S_{0},S_{1},\dots,S_{\mathrm{date}}\right).$ (8)
The process for updating long-term memory is defined as follows:
$Z^{\prime}_{\mathrm{date}+1}=S_{\mathrm{date}}\cup Z^{\prime}_{\mathrm{date}},$ (9)
$\mathrm{date}\in\{0,1,\dots,\mathrm{DAYS}\}.$
where $Z^{\prime}_{\mathrm{date}}$ represents the long-term memory, and $\mathrm{DAYS}$ denotes the maximum number of days.
Together, short-term and long-term memory provide essential context for the agents. Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors. Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
Figure 8: Overview of Complete Workflow.
Appendix B NASDAQ Stock
This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA. These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes. The NASDAQ stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. This dataset reflects current market trends and serves as a timely foundation for our research. It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
The training and testing periods for MACD, StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and our system are shown in Table 7. StockFormer and TimesNet require longer training periods, with their training datasets spanning a broader time range compared to the other models. In contrast, our system does not require training and relies solely on historical stock data for making trading decisions.
Strategy Training Testing
Start End Start End
MACD 29/7/2024 30/8/2024 3/9/2024 29/10/2024
StockFormer 3/7/2023 30/8/2024 3/9/2024 29/10/2024
TimesNet 3/7/2023 30/8/2024 3/9/2024 29/10/2024
Ours 16/8/2024 30/8/2024 3/9/2024 29/10/2024
Table 7: Training and Testing Periods for Various Models on the NASDAQ Stock Dataset.
The testing period is set from September 3, 2024, to October 29, 2024, to prevent potential data leaks that could provide prior knowledge to the GPT-4o system. This timeframe ensures the fairness of the evaluation by mitigating biases from such leaks, thereby enhancing the reliability of the experimental outcomes.
Appendix C Visualization Input
Figure 7 illustrates the system’s input prompts and corresponding outputs during the strategy update process. The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
Appendix D Simulation Process
In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis. Next, historical stock data is analyzed, followed by decision-making and execution. Short-term memory is formed through interactions with the environment. Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory. This entire process is illustrated in Figure 8.
Appendix E Simulation Process
The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), DeepSeek (Liu et al., 2024), Qwen-2.5 (Yang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, DeepSeek-chat, Qwen2.5-72B-Instruct, and Gemini-1.5-pro.

### v2

@@
-LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena
+Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents
 Abstract
-Recent advancements in large language models (LLMs) have significantly improved performance in natural language processing tasks.
-However, their ability to generalize to dynamic, unseen tasks, particularly in numerical reasoning, remains a challenge.
-Existing benchmarks mainly evaluate LLMs on problems with predefined optimal solutions, which may not align with real-world scenarios where clear answers are absent.
-To bridge this gap, we design the Agent Trading Arena, a virtual numerical game simulating complex economic systems through zero-sum games, where agents invest in stock portfolios.
-Our experiments reveal that LLMs, including GPT-4o, struggle with algebraic reasoning when dealing with plain-text stock data, often focusing on local details rather than global trends.
-In contrast, LLMs perform significantly better with geometric reasoning when presented with visual data, such as scatter plots or K-line charts, suggesting that visual representations enhance numerical reasoning.
-This capability is further improved by incorporating the reflection module, which aids in the analysis and interpretation of complex data.
-We validate our findings on NASDAQ Stock dataset, where LLMs demonstrate stronger reasoning with visual data compared to text.
-Our code and data are publicly available at https://github.com/wekjsdvnm/Agent-Trading-Arena.git.
-LLM Knows Geometry Better than Algebra:
-Numerical Understanding of LLM-Based Agents in A Trading Arena
-Tianmi Ma1, Jiawei Du3,4,, Wenxin Huang2, ✉, Wenjie Wang5, Liang Xie6, Xian Zhong1, ✉, and Joey Tianyi Zhou3,4 1 Hubei Key Laboratory of Transportation Internet of Things, Wuhan University of Technology 2 Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University 3 Centre for Frontier AI Research, Agency for Science, Technology and Research, Singapore 4 Institute of High Performance Computing, Agency for Science, Technology and Research, Singapore 5 School of Computing, National University of Singapore  6 School of Science, Wuhan University of Technology
-footnotetext: Email:wenxinhuang_wh@163.com, zhongx@whut.edu.cn. denotes the equal contribution. ✉ represents the corresponding author.
+Large language models (LLMs) have demonstrated remarkable capabilities in natural language tasks, yet their performance in dynamic, real-world financial environments remains underexplored.
+Existing approaches are limited to historical backtesting, where trading actions cannot influence market prices and agents train only on static data.
+To address this limitation, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive multi-agent trading and directly impact price dynamics.
+By simulating realistic bid-ask interactions, our platform enables training in scenarios that closely mirror live markets, thereby narrowing the gap between training and evaluation.
+Experiments reveal that LLMs struggle with numerical reasoning when given plain-text data, often overfitting to local patterns and recent values.
+In contrast, chart-based visualizations significantly enhance both numerical reasoning and trading performance.
+Furthermore, incorporating a reflection module yields additional improvements, especially with visual inputs.
+Evaluations on NASDAQ and CSI datasets demonstrate the superiority of our method, particularly under high volatility.
+All code and data are available at https://github.com/wekjsdvnm/Agent-Trading-Arena.
+Agent Trading Arena: A Study on Numerical Understanding in LLM-Based Agents
+Tianmi Ma1,#\#, Jiawei Du2,#\#, Wenxin Huang3, ✉, Wenjie Wang4, Liang Xie1, Xian Zhong1, ✉, and Joey Tianyi Zhou2 1 Wuhan University of Technology  2 Agency for Science, Technology and Research, Singapore 3 Hubei University  4 University of Science and Technology of China Correspondence: wenxinhuang_wh@163.com, zhongx@whut.edu.cn
+footnotetext: #\# denotes equal contribution.
 1 Introduction
-Recent advancements in large language models (LLMs) have demonstrated exceptional proficiency across various domains, achieving state-of-the-art performance in natural language processing (NLP) tasks such as translation (Koshkin et al., 2024), summarization (Yang et al., 2023), and reasoning (Kalyanpur et al., 2024; Qiao et al., 2023).
-While LLMs excel in language-based tasks, further progress in numerical and geometric reasoning is essential to tackling complex, interdisciplinary challenges, particularly in fields like finance and scientific research.
-Benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been developed to assess and improve the mathematical problem-solving abilities of LLMs through structured datasets and standardized evaluation protocols.
-These benchmarks not only advance LLM development but also represent a crucial step in enhancing the ability of artificial intelligence systems to solve mathematical reasoning problems.
-Figure 1: Illustration and Performance Comparison of Textual and Visual (K-Line Chart) Numerical Inputs.
-Top: The left figure presents numerical data in K-line chart format, while the right figure shows textual numerical inputs.
-Bottom: LLMs with visual inputs (orange) significantly outperform those with textual inputs (blue) in 40-day yield trends, demonstrating the advantage of visual data in financial decision-making.
-In response, recent updates to LLMs have focused on improving their mathematical abilities, with models (Ahn et al., 2024; Romera-Paredes et al., 2024) achieving new records on math-focused benchmarks.
-However, many of these benchmarks (Taylor et al., 2022; Naveed et al., 2023) primarily evaluate performance on problems familiar to the models, closely resembling those encountered during training and often following standard formats, solution strategies, or recurring mathematical patterns.
-This raises concerns about whether LLMs genuinely possess reasoning abilities for novel numerical data or whether their success is largely due to memorization and pattern recognition.
-These limitations underscore the need for alternative evaluation paradigms that assess LLMs’ ability to generalize mathematical principles to unseen scenarios.
-To address these limitations, we introduce the Agent Trading Arena, a virtual numerical game generating diverse numerical data through agent interactions in a zero-sum stock market environment.
-In this system, LLM-based agents make trading decisions based on historical stock prices, dynamically determined by agents’ bidding activities.
-Unlike static benchmarks (Naveed et al., 2023) with predefined strategies, our simulation forces agents to adapt to evolving market conditions.
-Any temporarily optimal strategy is quickly countered, fostering an environment that rewards adaptability.
-By mimicking competitive trading scenarios where outcomes depend on real-time agent decisions, we create a controlled yet dynamic environment that tests LLMs’ ability to infer hidden patterns and generalize numerical laws without relying on memorization or prior knowledge.
-Our experiments reveal that LLMs struggle with textual numerical data, often focusing on absolute values rather than capturing percentage changes and relations between data points, akin to algebraic reasoning.
-Additionally, LLMs tend to overemphasize recent data while overlooking earlier information, even when explicitly highlighted in the prompt.
-These limitations suggest that LLMs lack the ability to abstract numerical information into higher-level representations, a crucial skill for generalizing beyond explicit values.
-In contrast, LLMs perform better when presented with numerical data in visualized formats, such as scatter plots, line charts, and bar graphs, which involve geometric reasoning.
-Our experiments show that LLMs processing visual numerical data consistently outperform those processing textual numerical data.
-As shown in Figure 1, visual numerical data enables LLMs to achieve significantly higher return rates in the Agent Trading Arena, underscoring the advantages of structured visual representations.
-Incorporating the reflection module (Zhou et al., 2022; Shinn et al., 2023) further emphasizes the performance gap between textual and visual data, with the module proving particularly effective for visual representations.
-By leveraging structured visualizations, the reflection module enhances the model’s ability to reason through complex data relations, leading to more accurate and strategic decision-making.
-The Agent Trading Arena provides a valuable framework for advancing LLM generalization and adaptability, with broad applications in fields like finance, healthcare, and scientific research.
-To further validate our findings, we extend our evaluation to real-world financial data, such as NASDAQ Stock dataset, which includes historical trading data.
-The consistent results suggest an important contribution: LLMs exhibit stronger reasoning with visual geometric data compared to textual numerical data, demonstrating the value of visual representations in enhancing LLM performance.
+In recent years, large language models (LLMs) have demonstrated strong capabilities in contextual understanding and generation across tasks such as natural language processing (Koshkin et al., 2024), data analysis, and reasoning (Olausson et al., 2023; Qiao et al., 2023).
+Their applications are rapidly extending to complex decision-making and interactive scenarios.
+LLM-based autonomous agents have emerged as a promising paradigm, showing notable progress in communication (Qian et al., 2024; Zhong et al., 2024), social simulation (Park et al., 2023, 2022; Lin et al., 2023), gaming (Wang et al., 2024a; Zhu et al., 2023), and multi-agent collaboration (Chen et al., 2023; Hu et al., 2023; Hong et al., 2024).
+These agents are increasingly regarded as a pathway toward general artificial intelligence, with growing potential in dynamic decision-making and interactive environments.
+Figure 1: Comparison between Static Backtesting and the Arena.
+Left: Static backtesting performs one-shot analysis on historical data with fixed trading decisions, leveraging models such as Agent (market reasoning), StockFormer (predictive coding + RL hybrid), and StockMixer (time-series prediction).
+Right: The Arena, by contrast, supports real-time sensing, continuous adaptation, and feedback-driven execution for dynamic trading.
+Despite these advances, the application of LLM-based agents in financial markets remains limited.
+Financial ecosystems naturally form complex multi-agent environments in which participants interact through dynamic market mechanisms.
+However, most existing studies (Cheng and Chin, 2024; Xiao et al., 2024) rely on historical backtesting, as illustrated in Fig.˜1, where agents cannot influence market prices and thus fail to capture interactivity and adaptive behavior.
+This gap underscores the need for a framework that supports real-time agent interactions and realistic market feedback.
+To address these limitations, we present the Agent Trading Arena, a virtual zero-sum stock market in which LLM-based agents engage in competitive trading and directly affect price dynamics through bid-ask interactions.
+Our platform generates continuously evolving numerical data, forcing agents to adapt to shifting market conditions; any temporarily optimal strategy is quickly countered, rewarding adaptability and robust reasoning over memorization.
+By simulating bid-ask order matching in real time, the Arena faithfully reproduces liquidity dynamics and price impact, enabling agents to experience authentic market friction and slippage.
+Unlike static backtests, the Arena continuously updates its state based on agents’ collective actions, creating an ever-shifting landscape that prevents overfitting to historical patterns.
+Because agents directly influence subsequent price trajectories, the Arena closes the loop between training and evaluation, effectively bridging the “train-test” gap of traditional financial simulations.
+Its modular design further supports seamless integration of new market features, such as varying transaction costs, order types, or macroeconomic shocks, making it a versatile testbed for financial AI research.
+Our experiments show that LLMs struggle with numerical reasoning when processing plain-text data: they focus on absolute values, overlook percentage changes and relational patterns, and overemphasize recent trends even when earlier data are emphasized.
+In contrast, visual representations, such as line charts, bar graphs, and scatter plots, substantially improve both numerical reasoning and trading performance.
+As shown in Table˜5, chart-based inputs lead to significantly higher return rates in the Agent Trading Arena.
+We further integrate a reflection module (Zhou et al., 2022; Shinn et al., 2023), which amplifies the benefits of visual data by enabling agents to iteratively refine their reasoning.
+This module is particularly effective with structured visual inputs, facilitating deeper pattern recognition and more strategic decision-making.
+Finally, we validate our approach on real-world datasets from NASDAQ and CSI, demonstrating consistent performance gains over baselines, particularly under high-volatility market regimes.
+Our contributions are summarized threefold:
+  • We design the Agent Trading Arena, a dynamic and interactive simulation framework that moves beyond static backtesting with zero-sum multi-agent competition, enabling realistic evaluation of LLMs’ reasoning and adaptability.
+  • We uncover LLMs’ limitations in textual numerical reasoning, highlighting difficulties with relational patterns, percentage changes, and long-range dependencies, while demonstrating superior performance with structured visual data.
+  • We demonstrate that a reflection module enhances LLMs’ reasoning over complex visual inputs, leading to more accurate and strategic trading decisions.
 2 Related Works
-2.1 Mathematical Benchmarks for LLMs
-Math word problems (MWPs) have been widely studied, leading to the development of various benchmarks for evaluating models’ mathematical reasoning and problem-solving abilities.
-Early datasets, such as MAWPS (Koncel-Kedziorski et al., 2016), standardized existing problems to facilitate consistent evaluation.
-Math23K (Wang et al., 2017) introduced a large-scale collection of Chinese arithmetic problems that require structured equation solving.
-To increase diversity, benchmarks like ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021) provide richer annotations and a broader range of problem types.
-More recent benchmarks, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), focus on multi-step reasoning and advanced mathematical concepts, broadening the scope of evaluation.
-Additionally, MathQA-Python (Austin et al., 2021), a Python variant of MathQA (Amini et al., 2019), emphasizes programmatic reasoning, while MGSM (Shi et al., 2023) extends these benchmarks to multilingual contexts.
-Despite these advancements, current models (Lu et al., 2023b, a) primarily rely on memory-based answering strategies learned through extensive training, rather than demonstrating true mathematical reasoning.
-2.2 LLMs for Enhanced Mathematical Reasoning
-Building on these benchmarks, LLMs have advanced mathematical problem-solving by incorporating specialized datasets into their training.
-Models such as Galactica (Taylor et al., 2022), PaLM-2 (Anil et al., 2023), Minerva (Lewkowycz et al., 2022), and LLaMA-2 (Touvron et al., 2023) leverage extensive datasets during pre-training, improving their mathematical reasoning abilities and understanding of complex concepts.
-Fine-tuned models like MetaMath (Yu et al., 2024), MAmmoTH (Yue et al., 2024), and WizardMath (Luo et al., 2023) are specifically tailored for mathematical tasks.
-These models undergo domain-specific fine-tuning with carefully curated datasets, enabling them to tackle advanced reasoning challenges and significantly enhance performance on specialized benchmarks.
-However, existing methods (Naveed et al., 2023; Yang et al., 2024a) often rely on large training datasets, raising concerns about the true reasoning capabilities of LLMs.
-While these methods primarily assess performance on established problem types, the heavy reliance on extensive training data suggests that models may achieve high performance through memorization and pattern recognition, rather than genuine reasoning.
-Consequently, alternative evaluation paradigms are needed to more accurately assess LLMs’ ability to generalize mathematical principles to novel scenarios.
-Figure 2: Stock Trading Workflow in Agent Trading Arena.
-Top: Workflow of a trading day, including preparation, trading, and post-trading reflection.
-Agents discuss insights in the chat pool, analyze market trends, execute trades, and refine strategies based on performance.
-Bottom: Example of agents’ interactions in the chat pool and dynamic strategy updates.
+2.1 LLM-Based Agents in Dynamic Decision-Making Environments
+LLMs serve as the core reasoning engines in autonomous agents, enabling planning, tool use, reflection, and adaptation within dynamic environments.
+Advances in NLP have empowered models such as ChatGPT (OpenAI, 2023), PaLM (Anil et al., 2023), and LLaMA (Touvron et al., 2023) to operate via natural language, making them particularly effective in interactive, text-based scenarios.
+Frameworks like ChatDev (Qian et al., 2024), AgentVerse (Chen et al., 2023), and MetaGPT (Hong et al., 2024) demonstrate multi-agent collaboration on software development tasks, while Voyager (Wang et al., 2024a), GROOT (Cai et al., 2024), and Optimus-1 (Li et al., 2024b) showcase autonomous exploration in open-ended environments such as Minecraft.
+Social simulators, including Generative Agents (Park et al., 2023) and AgentSims (Lin et al., 2023), further reveal emergent behaviors through sustained multi-agent interactions.
+Together, these studies demonstrate that LLM-based agents can effectively coordinate, communicate, and make decisions in complex, evolving environments.
+Their adaptability and collaborative strengths highlight strong potential for real-time decision-making in finance, an information-dense and competitive domain analogous to the dynamic settings where LLMs excel.
+2.2 LLMs in Financial Analysis and Decision Making
+LLMs have been applied to sentiment analysis, financial question answering, and investment recommendation.
+Benchmarks such as SocioDojo (Cheng and Chin, 2024) evaluate financial understanding but lack interactive components.
+Multi-agent systems like StockAgent (Zhang et al., 2024a), TradingAgents (Xiao et al., 2024), and CryptoTrade (Li et al., 2024a) enhance performance through role specialization and iterative reflection.
+In contrast, traditional forecasting models, e.g., StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), PDFormer (Jiang et al., 2023), StockMixer (Fan and Shen, 2024), and iTransformer (Liu et al., 2024b), focus on price prediction using advanced methods such as wavelet transforms and temporal modeling to capture market dynamics.
+However, neither paradigm addresses the “live market” paradox: without real-time adaptive mechanisms, these systems cannot simulate how participants adjust strategies in response to multi-agent interactions, news shocks, and evolving order flows.
+This limitation significantly reduces their practical utility in real trading environments.
+Figure 2: Overall architecture of the Agent Trading Arena, consisting of three components: the Chat Pool (left) for inter-agent communication, agent modules (center) for analysis, decision-making, and reflection, and the trading platform (right), where agents’ buy/sell actions dynamically determine asset prices in a closed-loop market simulation.
+3 Agent Trading Arena
+To overcome the limitations of static backtesting and capture dynamic interactions among market participants, we propose the Agent Trading Arena, a closed-loop, zero-sum economic system that simulates complex quantitative scenarios in real markets (Guo et al., 2024).
+The overall architecture is shown in Fig.˜2 and detailed in Appendix˜A.
+Within this environment, agents invest in assets, earn dividends on their holdings, and pay daily expenses using virtual currency.
+3.1 Virtual Market Environment
+Closed-Loop System.
+The Agent Trading Arena is a closed-loop virtual economy that isolates external knowledge and avoids reliance on historical data, enabling fair evaluation of agent capabilities.
+Unlike traditional backtesting grounded in past market records, this sandboxed system removes real-world anchors: all assets are abstract symbols, and their prices emerge solely from agent interactions through a bid-ask mechanism.
+By eliminating external context, the environment compels agents to adapt strategies based on internal dynamics and opponent behavior, fostering genuine decision-making and strategic evolution.
+Bid-Ask Mechanism.
+Asset prices arise exclusively from a bid-ask process: each agent’s orders directly influence the market price, with no external benchmarks.
+This closed-loop design rigorously captures endogenous price formation, faithfully modeling liquidity constraints, order book dynamics, market depth, and slippage.
+Every execution reshapes the market microstructure, producing non-trivial price impacts and path-dependent outcomes that evolve with aggregate agent behavior over time.
+3.2 Chat Pool
+To simulate real-world information flow, we introduce a shared Chat Pool where agents post and read “news” items, analyst commentary, and social media-style opinions.
+By reproducing both informative signals and noisy interference, including hallucinations (Gekhman et al., 2024) and misinformation, the Chat Pool reflects how public information can shape, and at times distort, traders’ narratives and expectations.
+This process increases the complexity of strategic reasoning and adaptive behavior in the market.
+3.3 Market Mechanisms
+We incorporate two key incentives to sustain active trading and realistic dynamics: 1) Dividends & Capital Gains: Agents earn income from both price appreciation and dividends, distributed proportionally to holdings.
+This implicitly anchors prices and rewards low-cost positions.
+2) Daily Wealth Fee: A cost proportional to each agent’s total wealth is charged daily.
+Only those earning sufficient dividends can cover this fee, which discourages passive holding and promotes frequent, strategic trades.
+Together, these mechanisms maintain continuous market activity, encourage adaptive decision-making, and prevent reliance on static or exploitative strategies in a zero-sum environment.
+3.4 Zero-Sum Competition
+By design, the Arena is zero-sum: the profit of one agent is the loss of another, meaning there is no fixed “best” strategy.
+Competitors counter any temporarily successful tactic, compelling agents to learn implicit rules, interpret delayed numerical feedback, and continuously adapt in real time.
+With results unfolding gradually and misinformation potentially introduced by rivals, agents must rely on experiential learning and develop context-aware strategies through ongoing adaptation rather than predefined rules or external knowledge.
 4 Proposed Method
-To mitigate the influence of human prior knowledge and memory, we designed a closed-loop economic system (Guo et al., 2024) called the Agent Trading Arena, a zero-sum game simulating complex, quantitative real-world scenarios.
-The simulation workflow is illustrated in Figure 2 and further detailed in Appendix A.
-In the Agent Trading Arena, agents can invest in assets, earn dividends from holding assets, and pay daily expenses using virtual currency.
-The agent with the highest total return wins the game.
-3.1 Agent Trading Arena
-Structure of Agent Trading Arena.
-To eliminate external knowledge biases, asset prices are determined by a bid-ask system, reflecting the prices at which buyers and sellers are willing to transact.
-The system evolves solely based on agents’ actions and interactions, without external influences.
-This design ensures that the outcomes of agents’ actions are not immediately apparent but unfold gradually, influenced by other agents’ decisions.
-To encourage active participation, a dividend mechanism is introduced.
-There are two primary sources of income in this system: capital gains from asset price differentials and dividends from holding assets.
-Dividends for each asset are distributed according to a predefined ratio, serving as an implicit anchor for asset prices.
-Agents holding more low-cost assets receive higher dividends.
-To prevent passive asset holding until the end of the game, agents must pay a daily capital cost proportional to their total wealth.
-These expenses are offset by asset dividends, and only agents with sufficient low-cost assets can cover costs.
-Under the pressure of significant daily expenses, agents must act swiftly and strategically, triggering frequent trades and price fluctuations to stimulate market activity.
-This dynamic mechanism ensures fairness in the zero-sum game while preventing agents from relying on fixed strategies to find optimal solutions.
-Agents Learn and Compete in Arena.
-The zero-sum game structure is crucial to eliminating the possibility of a universally optimal strategy.
-In fixed scenarios with a static optimal solution, agents could rely on predefined rules or memory-based approaches, bypassing adaptive decision-making.
-The zero-sum game ensures that there is no universally correct solution, with outcomes evolving dynamically based on agent interactions and competition.
-This design forces agents to continually adapt, learn from feedback, and develop context-dependent strategies, promoting deeper environmental exploration and preventing reliance on static or memory-driven solutions.
-In the Agent Trading Arena, agents are unaware of implicit rules, except for the objective to maximize their virtual wealth throughout the simulation.
-To win this zero-sum game, agents must effectively learn from experience, decipher hidden game rules, and develop strategies to counter competitors.
-This requires the ability to comprehend numerical feedback, formulate enduring strategies, and make informed decisions.
-Unlike other mathematical reasoning problems, the results of their actions unfold gradually and dynamically.
-Moreover, agents are easily misled by erroneous information from competitors, hindering their ability to discern strategic cues from competitors’ textual data.
-Importantly, agents remain unaware of these implicit rules, so applying real-world knowledge does not benefit their performance.
-Therefore, agents must rely on experiential learning to decipher the hidden game rules and ultimately achieve victory.
-3.2 Types of Numerical Data Input
-Limitations of Textual Numerical Data.
-In the Agent Trading Arena, the generated stock data is stored in numerical format.
-When used directly as input to an LLM, the models often struggle to interpret numerical data accurately or make sound decisions.
-To mitigate this, we convert the data into textual formats (Hu et al., 2024; Zhang et al., 2024), enhancing semantic features and clarifying output requirements to improve the models’ understanding.
-During interactions, the LLMs process stock prices, trading volumes, and market indices presented as textual numerical data.
-Figure 3: Textual and Visual Representations of Corresponding Inputs and Outputs.
-The left images display the agent’s Buy and Sell trading records, daily trade prices, and K-line charts for three stocks.
-The output from visual inputs (bottom right) captures overall stock trends and long-term behavior, while the output from textual inputs (top right) focuses on specific current prices.
-However, this textual approach reveals significant limitations.
-While the data is presented clearly, LLMs tend to focus excessively on specific values rather than identifying long-term trends or global patterns.
-They also struggle with understanding correlative relations and percentage changes, limiting their ability to assess differences and identify connections between data points.
-When analyzing time-series data with complex patterns, LLMs often fixate on individual data points, overlooking overarching relations.
-This issue is evident in the analysis output in the top-right corner of Figure 3, where LLMs’ focus on individual values impedes their ability to generalize, reducing their capacity to extract meaningful global insights.
-Additionally, LLMs often overemphasize recent data while undervaluing historical information, even when prompted to consider its importance.
-This prevents them from effectively integrating past data and recognizing long-term patterns, complicating their understanding of numerical relations and trends.
-These challenges highlight the need for improved mechanisms to process numerical relations, identify global trends, and derive deeper insights from textual numerical data.
-Potential of Visual Numerical Data.
-Since textual numerical data often leads LLMs to focus on local details while neglecting broader relations, we investigated whether visual representations, such as scatter plots, line charts, and bar charts, could help LLMs better understand overall trends, similar to human reasoning.
-Thus, we transition from textual numerical data inputs to visualized formats  (Yang et al., 2025).
-As demonstrated in the bottom-right corner of Figure 3, visual representations enable LLMs to more effectively grasp global trends, patterns, and relations that are often difficult to discern from textual numerical data alone.
-These findings highlight the advantages of structured, visual numerical data, indicating that this format allows LLMs to more intuitively and comprehensively understand complex data, better capturing overall fluctuations, whereas text tends to focus on local details.
-By combining visualization and textual representations, LLMs not only overcome the challenges of relations in time-series data but also demonstrate better performance in identifying long-term trends and global patterns, while still attending to local details.
-3.3 Reflection Module
-We propose a strategy distillation method, illustrated in Figure 4, that delivers real-time feedback to LLMs by analyzing both descriptive textual and visual numerical data.
-This enables the generation of new strategies and optimization of action plans.
-The approach allows agents to evaluate their results, refine strategies, and adapt continuously based on feedback.
-The process begins with assessing the day’s trajectory memory and associated strategies using an evaluation function.
-The strategic generation process leverages contrastive analysis of peak and nadir performers from the evaluation phase, creating bidirectional learning signals that inform subsequent iterations.
-This iterative cycle ensures continuous strategy evolution, fostering sustained improvement in decision-making.
+Our agent, ArenaTrader, integrates textual and visual modalities to form a comprehensive understanding of market conditions.
+It employs a structured three-stage framework, market analysis, trade execution, and strategic reflection, to leverage this multi-modal representation.
+This design enables informed and adaptive decision-making based on data-driven insights and iterative feedback.
+Figure 3: Textual vs.
+Visual Input–Output Representations.
+Left: Agent buy/sell records, daily trade prices, and corresponding candlestick charts.
+Top right: Visual-input outputs highlight overall trends and long-term patterns.
+Bottom right: Textual-input outputs emphasize specific price points.
+4.1 Numerical Data Representations
+Textual Input.
+Time-series stock data are represented as plain text, including timestamps, prices, and volumes, to provide explicit semantics (Hu et al., 2024; Zhang et al., 2024b).
+However, LLMs often fixate on absolute values, overlook percentage changes, and overemphasize recent entries, which limits their ability to reason across long horizons (see lower right of Fig.˜3).
+Visual Input.
+To complement textual data, we introduce visualizations such as line charts, bar graphs, and scatter plots (Yang et al., 2025), which highlight global trends, relational patterns, and volatility at a glance (see upper right of Fig.˜3).
+These visualizations are consistent with the textual input, ensuring a cohesive representation.
+This structured format enables LLMs to integrate local details with broader market context more effectively.
+4.2 Agent Architecture
+Investment Analysis.
+Inspired by human trading behavior, each agent begins by analyzing market conditions before executing trades.
+This process involves extracting and synthesizing multi-modal signals from the Chat Pool, price histories, and fundamental indicators.
+By combining these heterogeneous inputs, the agent constructs a stock ranking framework based on predefined heuristics or learned criteria, allowing it to identify promising assets for portfolio allocation.
+Action Decision.
+Agents execute multiple trades per day, making buy, sell, or hold decisions by integrating diverse signals, including analysis reports, real-time market data, and historical performance trends.
+They evaluate volatility, price momentum, and portfolio risk exposure to compute optimal trade volumes under capital constraints.
+Orders are executed via the bid-ask mechanism, with portfolios updated accordingly.
+Each decision aims to maximize long-term returns while maintaining a competitive edge in the evolving virtual market.
 Figure 4: Design of the Reflection Module.
-The process evaluates daily trajectory memory and strategies (top right), then generates new strategies (center) based on evaluation, environmental feedback (bottom right), and feedback from the 5 top- and bottom-performing strategies.
-Stock visualization (bottom left) enhances reflection, driving continuous improvement.
-The reflection module plays a crucial role in refining strategies by offering real-time feedback.
-It analyzes both descriptive textual and visual numerical data to generate new strategies and optimize action plans.
-Within the Agent Trading Arena, the reflection module is triggered regularly to consolidate daily trading records and evaluate the effectiveness of strategies, refining both successful and unsuccessful experiences to guide future decisions.
-Ineffective strategies are stored in a strategy library for future reference, allowing agents to review and learn from past experiences.
-Further details can be found in Appendix A.
-ID Name Year Cash Occupation
+Left: Agent trading records, strategy evaluations, and top- and bottom-performing strategies from the strategy library.
+Top right: Corresponding stock chart.
+Bottom right: Newly generated strategy based on current information.
+Strategy Optimization via Reflection.
+At the end of each trading day, the reflection module distills daily experiences into bidirectional learning signals (see Fig.˜4).
+It first evaluates the effectiveness of trading trajectories and strategies using a scoring function.
+Then, by contrasting top- and bottom-performing strategies, agents generate refined action plans and archive unsuccessful tactics in a strategy library.
+This iterative process drives continuous strategy evolution and more robust decision-making.
+5 Experimental Results
+5.1 Experimental Setup
+Environment.
+The Agent Trading Arena is lightweight and accessible, requiring only API credits and no GPU for deployment. Details on the selection of different LLMs are provided in Appendix˜D.
+ID Name
+Duration
+Initial
+Capital
+Profession
 1 Amy 1 100,000 AI Researcher
@@
 9 Ivy 3 100,000 Marketing Manager
-$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
-Table 1: Agent Details. “Year” is the investment duration, and “Cash” is the initial capital.
-ID Name DPS Historical Close Prices Quantity
-1 A 22 454.17, $\dots$, 445.60 1,200
-2 B 23 354.17, $\dots$, 465.80 1,000
-3 C 25 500.47, $\dots$, 440.60 1,600
-$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
-Table 2: Stock Details. “DPS” is the Dividend Per Share, and “Quantity” is the initial share count.
-4 Experimental Results
-4.1 Experimental Setup
+Table 1: Agent Characteristics. Duration denotes each agent’s investment horizon, and Initial Capital indicates the starting funds for trading.
+ID Ticker DPS Closing Prices Initial Quantity
+1 A 22 454.17, …, 445.60 1,200
+2 B 23 354.17, …, 465.80 1,000
+3 C 25 500.47, …, 440.60 1,600
+Table 2: Stock Details. DPS denotes dividend per share; Closing Prices are historical end-of-day closing prices; and Initial Quantity specifies the starting number of shares.
 Datasets.
-To evaluate the ability of LLMs to analyze and process data, we developed the Agent Trading Arena, a controlled environment that isolates external factors.
-The system’s workflow is illustrated in Figure 2.
-Within this environment, agents discuss stock market trends, analyze stock data, and engage in trading activities.
-Each agent can execute multiple trades per day and reflect on all trades at the end of each trading session.The Agent Trading Arena allows for adjustable numbers of agents and stocks; in our experiment, we deployed at least nine agents and three stocks.
-All agents were provided with the same initial capital to ensure identical starting conditions.
-Detailed information about each agent is presented in Table 1, and stock-specific data is available in Table 2.
-To further validate our findings, we selected a subset of NASDAQ Stock dataset for portfolio investment.
-For more details, refer to Appendix B.
-Evaluation Metrics.
-Each agent was assigned varying capital based on their roles, and performance was evaluated using the following metrics:
-1) Total Return (TR): Measures the overall performance of the strategy, calculated as: $\mathrm{TR}={(C_{1}-C_{0})}/{C_{0}}$, where $C_{0}$ is the initial asset value and $C_{1}$ is the final asset value.
-2) Win Rate (WR): Represents the proportion of winning trades, calculated as: $\mathrm{WR}={N_{w}}/{N_{t}}$, where $N_{w}$ is the number of winning trades and $N_{t}$ is the total number of trades.
-3) Sharpe Ratio (SR): Evaluates investment returns relative to risk, calculated as: $\mathrm{SR}={(R_{p}-R_{f})}/{\sigma_{p}}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is the risk-free return, set to 0 as defined in SocioDojo (Cheng and Chin, 2024).
-4) Mean Daily Return (Mean): Indicates the average return per day during the trading period.
-5) Standard Deviation of Daily Returns (Std): Reflects the volatility of daily returns, indicating the risk associated with the strategy.
-4.2 Comparative Experiments
-We conducted experiments to assess the real-time data analysis and reasoning capabilities of LLM-based agents, focusing on how textual and visual representations influence decision-making.
-First, we explored the impact of textual and visual representations in dynamic environments.
-Next, we incorporated the reflection module to enhance agents’ reasoning and data interpretation, examining how reflective reasoning influences decision-making.
-Finally, we validated the model’s effectiveness through stock investment simulations on NASDAQ Stock dataset, assessing agents’ adaptability and decision-making in real-world scenarios.
-LLMs Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
-w/o Reflection LLaMa-3 (Dubey et al., 2024) 4.0934 0.6857 2.2700 62.5001 0.1963
-DeepSeek (Liu et al., 2024) 27.3078 2.2478 2.7553 92.2078 1.5303
-Qwen-2.5 (Yang et al., 2024b) 30.3740 2.6795 1.4588 93.7500 1.7025
-Gemini-1.5 (Reid et al., 2024) 14.3193 1.2210 2.7776 83.8384 1.7293
-19.0389 1.6038 1.1654 90.9091 1.3761
-23.9649 1.9809 1.3342 100.0000 1.4847
-GPT-4o (Hurst et al., 2024) 13.0369 1.1386 1.9285 54.5455 0.5904
-17.0661 1.4539 1.5040 72.7272 0.9668
-26.1806 2.1574 2.0578 90.9091 1.0484
-w/ Reflection LLaMa-3 (Dubey et al., 2024) 10.2458 1.6669 2.4034 66.6667 0.6936
-DeepSeek (Liu et al., 2024) 30.6238 2.4696 1.5121 100.0000 1.6332
-Qwen-2.5 (Yang et al., 2024b) 38.9113 3.3244 1.0726 100.0000 3.0994
-Gemini-1.5 (Reid et al., 2024) 29.4511 2.5185 3.0116 93.5064 2.2100
-37.0054 2.9111 1.2349 100.0000 2.3574
-41.3264 3.1946 0.1588 100.0000 20.1128
-GPT-4o (Hurst et al., 2024) 33.6508 2.6713 2.1711 98.7013 2.1417
-35.7622 2.8206 0.6782 100.0000 4.1590
-47.6851 3.6095 0.5327 100.0000 6.7765
-Table 3: Performance Comparison Using Textual and Visual Data w/o and w/ Reflection.
-Evaluation of agent portfolios show that visual approaches outperformed textual ones, with the best results achieved by combining both.
-LLMs demonstrated a significant advantage in TR and SR, with visual inputs yielding greater improvements than textual inputs.
-The best and second-best results are highlighted in bold and underlined.
-Trials with Textual or Visual Input.
-To enhance LLMs’ understanding of complex data in the Agent Trading Arena, we transitioned from textual numerical inputs to visualized formats, including scatter plots, line charts, and bar charts.
-Three types of visualizations were used: daily K-line charts, transaction histories, and quantities traded by each agent.
-The experimental setup for the visualization group is shown in Figure 3 and detailed in Appendix C.
-In the Arena, LLMs without image input capabilities received only text input and did not perform reflection.
-For image-enabled LLMs, the first agent received only visual input, the second received both textual and visual input, and the others received only textual input.
-None of the agents had reflection capabilities.
-For details on the selected LLMs, please refer to Appendix E.
-We conducted experiments across different LLMs and the results for Gemini-1.5 (Reid et al., 2024) and GPT-4o (Hurst et al., 2024) in Table 3 show that agents with visual numerical input outperformed those with textual input alone.
-Agents receiving both textual and visual input achieved the best performance.
-This combination enables agents to focus on local details while also understanding overall trends, leading to optimal performance.
-This suggests that LLMs exhibit better reasoning skills with visual geometric data rather than textual numerical data, highlighting a clear distinction in their strengths in geometry and algebra.
-Strategy TR $\uparrow$ SR $\uparrow$
-MACD (Chong and Ng, 2008) 7.18 0.173
-StockFormer (Gao et al., 2023) 9.05 0.073
-TimesNet (Wu et al., 2023) 11.74 0.203
-GPT-4o + Textual 8.69 0.167
-GPT-4o + Visual 9.91 0.195
-GPT-4o + Textual + Visual (Ours) 12.23 0.291
-Table 4: Reproduced Backtesting Performance on NASDAQ Stock Dataset. The agent with textual and visual input outperformed the NASDAQ-100 by 53.97% in SR during the same period.
-LLMs T V TR $\uparrow$ SR $\uparrow$
-LLaMa-3 100.00 100.00
-DeepSeek + 35.03 + 45.63
-Gemini-1.5 + 10.99 + 14.93
-+ 11.49 + 29.47
-GPT-4o + 10.39 + 28.29
-+ 17.18 + 40.77
-Table 5: Performance Comparison in Trading Decisions w/o Reflection, Competing Pairwise with LLaMa-3.
-DeepSeek may possess unique strengths or optimizations that allow it to better adapt to the task’s complexities when competing with other models.
-LLMs TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
-LLaMa-3 (Dubey et al., 2024) 100.0000 100.0000 100.0000 100.0000 100.0000
-Gemini-1.5 (Reid et al., 2024) + 3.8631 + 5.9249 + 12.0426 - 4.0010 + 2.2920
-DeepSeek (Liu et al., 2024) + 3.1317 + 3.1309 + 1.7169 + 9.8039 + 1.7945
-Qwen-2.5 (Yang et al., 2024b) + 7.4578 + 7.3907 - 2.4801 + 50.0000 + 8.0524
-GPT-4o (Hurst et al., 2024) + 17.7636 + 16.9041 + 2.4136 + 37.9310 + 14.2775
-Table 6: Performance Comparison in Trading Decisions w/ Reflection, Competing Pairwise with LLaMa-3 Using Textual Data.
-The results show that in competition with reflection using textual numerical data, GPT-4o and Qwen-2.5 outperformed other models.
-Trials with Reflection and Textual or Visual Input.
-Building upon the previous experiments, we incorporated the reflection module to investigate its impact on the experimental outcomes.
-By introducing this module, we aimed to determine whether reflective processing could improve strategies and adaptability in complex financial scenarios.
-Our experiments with diverse LLMs demonstrate that reflective agents consistently outperformed non-reflective agents in stock trading.
-As shown in Table 3, text-only LLMs with reflection modules achieved enhanced returns, with GPT-4o (Hurst et al., 2024) reaching 33.65% total return using textual input alone, which increased to 47.70% with visual data.
-Combining both input types resulted in a 41.74% higher return compared to using textual input alone.
-Compared to agents without the reflection module, including the reflection module significantly enhanced each agent’s stock investment performance.
-LLMs with the reflection module showed the greatest improvement in agents with visual inputs, compared to those with textual inputs, in both the Total Return and Sharpe Ratio.
-This indicates that incorporating reflection further widens the gap between LLMs’ geometric reasoning with visual data and algebraic reasoning with textual numerical data, highlighting their strengths in both areas.
-Simulation with NASDAQ Stock dataset.
-To further validate our findings, we conducted a two-month investment simulation on NASDAQ Stock dataset, starting with an initial capital of 100,000 units for portfolio investment.
-NASDAQ Stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays.
-Detailed information about the models used is provided in Appendix B.
-The results in Table 4 show that, despite StockFormer (Gao et al., 2023) and TimesNet (Wu et al., 2023) benefiting from longer training periods and larger datasets, they underperformed compared to our model.
-Notably, our model achieved superior results without additional training, relying solely on historical stock data for decision-making.
-The Sharpe ratios for Apple, NASDAQ-100, and S&P 500 during the same period were 0.097, 0.189, and 0.205, respectively.
-The agent with textual and visual input outperformed the NASDAQ-100 and S&P 500 in Sharpe Ratio by 53.97% and 41.95%, respectively.
-Moreover, only visual input outperformed textual input, further highlighting LLMs’ stronger geometric reasoning abilities.
-4.3 Ablation Experiments
-Impact of Modality on LLM Competitiveness.
-We employed a relative evaluation method for this experiment.
-The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline.
-This setup aimed to explore the impact of different agents and modalities on LLM performance.
-The results are shown in Table 5.
-The findings indicate that DeepSeek (Liu et al., 2024) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
-Competition Among Different LLMs with Reflection.
-To investigate the role of reflection in competition dynamics using textual numerical data between LLMs, we conducted an ablation study in the Agent Trading Arena.
-Using a relative evaluation approach, the first agent employed various LLMs, while eight LLaMa-3-based agents served as the baseline for a fair comparison.
-This setup effectively isolated the influence of each LLM on performance.
-As shown in Table 6, in competition with reflection, GPT-4o (Hurst et al., 2024) and Qwen-2.5 (Yang et al., 2024b) outperformed other models, consistent with the findings in Table 3.
+We configure the Arena with varying numbers of agents and stocks, all initialized with the same capital (see Tables˜1 and 2).
+To validate our agents, we simulate portfolio investments on subsets of NASDAQ and CSI datasets.
+For NASDAQ, we run a two-month simulation (September 3–October 29, 2024) with an initial capital of 100,000 units, excluding weekends and holidays.
+For CSI, we conduct a one-year simulation (January 2–December 31, 2024) with the same capital, also excluding non-trading days.
+Both datasets, collected from Yahoo Finance and Baostock, provide daily records of opening, closing, high, and low prices, trading volume, and associated technical indicators, offering a holistic and up-to-date depiction of market behavior.
+Additional details are given in Appendix˜B.
+Baselines.
+We compare ArenaTrader against the following strategies:
+1) Buy & Hold: A passive strategy that purchases assets initially and holds them throughout the period without further trading.
+2) SMA (Gencay, 1996): A moving average crossover strategy used to detect bullish and bearish momentum. We evaluate window sizes [5, 10, 15, 20] and select the best based on validation results.
+3) ZMR (Poterba and Summers, 1988): A zone-based reversal strategy that trades on deviations from preset bounds, with thresholds and holding periods tuned via validation.
+4) MACD (Chong and Ng, 2008): A momentum strategy based on MACD and signal line crossovers, where MACD is the 12-day EMA minus the 26-day EMA, and the signal line is the 9-day EMA.
+5) StockFormer (Gao et al., 2023): A hybrid model combining predictive coding and reinforcement learning to capture future dynamics and asset correlations.
+We follow the default setup from the original implementation.
+6) TimesNet (Wu et al., 2023): Transforms 1D time series into 2D representations to model multi-period temporal variations.
+We follow the original setup and select the look-back window from [5, 10, 15, 20] based on validation.
+7) StockMixer (Fan and Shen, 2024): A lightweight MLP that mixes indicators, time, and stocks for efficient modeling of market dynamics. We follow the settings specified in the original work.
+Metrics.
+We evaluate using five metrics:
+1) Total Return (TR): $(C_{1}-C_{0})/C_{0}$, where $C_{0}$ and $C_{1}$ are initial and final capital.
+2) Win Rate (WR): $N_{w}/N_{t}$, the ratio of profitable trading days to total trading days.
+3) Sharpe Ratio (SR): $(R_{p}-R_{f})/\sigma_{p}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is set to 0 to emphasize risk-adjusted returns.
+4) Mean Daily Return (Mean): Average daily return over the simulation.
+5) Return Volatility (Std): Standard deviation of daily returns, measuring risk.
+5.2 Experimental Results
+In the Agent Trading Arena, we configure multiple agents and LLM backbones to promote diversity.
+The primary agent, ArenaTrader, consistently uses our full-featured design, including visual perception and reflection, enabling focused analysis.
+We evaluate ArenaTrader on real-world datasets, with corresponding prompts detailed in Appendix˜C.
+Model Avg. Trend TR (%) $\uparrow$ Mean (%) $\uparrow$ Std (%) $\downarrow$ WR (%) $\uparrow$ SR $\uparrow$ $\Delta$ (%) $\uparrow$
+ArenaTrader w/ Qwen-VL-128k -0.4632 33.3018 2.6526 1.0150 100.00 2.6135 33.7650
+ArenaTrader w/ GPT-4o-mini 5.4722 38.6460 3.0249 1.4235 100.00 2.1249 33.1738
+ArenaTrader w/ Gemini-1.5 7.2743 41.3264 3.1946 0.1588 100.00 20.1128 34.0521
+ArenaTrader w/ Qwen-VL-32k -0.5216 35.8150 3.1213 1.6079 90.00 1.9412 36.3366
+ArenaTrader w/ GPT-4o 3.5881 47.6851 3.6095 0.5327 100.00 6.7765 44.0970
+Table 3: Investment Performance of ArenaTrader with Different LLMs.
+Avg.
+Trend denotes the average market trend, and $\Delta$ indicates the gain relative to this trend.
+GPT-4o achieves the highest returns and stability, demonstrating superior data-driven reasoning among vision-capable LLMs.
+Bold and underlined values denote the best and second-best results, respectively.
+Refer to caption Refer to caption
+(a) Gemini-1.5 (b) Qwen-VL-128k
+Refer to caption Refer to caption
+(c) GPT-4o-mini (d) GPT-4o
+Figure 5: ArenaTrader Performance vs.
+Market Volatility.
+Total return of ArenaTrader in the Agent Trading Arena, compared with the most volatile single stock and the average return across all stocks.
+This comparison highlights the agent’s robustness under varying market conditions.
+Vision-Enhanced Agent Performance.
+Tables˜3 and 5 report the returns of ArenaTrader under different vision-capable LLMs.
+GPT-4o (Hurst et al., 2024) achieves the highest TR, followed by Gemini-1.5 (Reid et al., 2024).
+For volatility-adjusted returns, GPT-4o again leads, with Qwen-VL-32k (Wang et al., 2024b) second.
+These results highlight GPT-4o’s strong data-driven reasoning and robustness across varying volatility regimes, underscoring its ability to integrate visual and numerical signals in dynamic markets.
+Strategy NASDAQ CSI
+TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$ TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
+Buy & Hold 6.06 0.155 0.916 61.54 0.169 0.73 0.004 0.403 43.15 0.010
+SMA 0.19 0.006 0.561 46.15 0.011 -3.45 -0.014 0.319 33.20 -0.044
+ZMR 1.04 0.030 0.814 41.03 0.037 1.24 0.006 0.361 41.91 0.016
+MACD 5.62 0.146 1.033 58.97 0.141 -0.67 -0.002 0.373 40.66 -0.006
+StockFormer (Gao et al., 2023) 11.23 0.281 0.121 56.41 0.231 14.83 0.636 1.128 41.08 0.056
+TimesNet (Wu et al., 2023) 11.41 0.283 1.035 56.41 0.273 32.93 0.139 2.053 48.55 0.068
+StockMixer (Fan and Shen, 2024) 14.64 0.357 1.173 56.41 0.305 43.00 0.153 0.907 27.80 0.168
+ArenaTrader w/ GPT-4o (Ours) 15.99 0.450 1.110 58.97 0.348 48.93 0.176 1.432 51.45 0.123
+Table 4: Strategy Performance on NASDAQ and CSI. Our method demonstrates strong robustness and generalization, achieving near-optimal performance without additional training.
+Evaluation on Real-World Data.
+Using GPT-4o as the backbone, Table˜4 presents results on NASDAQ and CSI.
+Training details for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and StockMixer (Fan and Shen, 2024) are provided in Appendix˜B.
+Despite extensive training, these models underperform our method, which uses a 10-day input window selected via ablation.
+The Sharpe Ratio (SR) of NASDAQ-100 benchmark during the same period is 0.189, while CSI 300 records 0.054.
+In contrast, our model achieves 0.348 on NASDAQ-100 (an improvement of 0.159) and 0.123 on CSI 300 (an improvement of 0.069).
+(a) NASDAQ
+(b) CSI
+Figure 6: Backtesting Results on Real-World Stock Datasets.
+Two-month backtest on NASDAQ and one-year backtest on CSI.
+The proposed method consistently delivers near-optimal performance and outperforms all baselines, particularly during periods of high market volatility.
+Building on these results, Fig.˜6 provides a detailed analysis of portfolio trajectories and market conditions, illustrating our agent’s near-optimal performance, especially during high-volatility periods.
+Notably, our approach operates without task-specific retraining, relying solely on historical price data.
+This design underscores the robustness and strong generalization ability of our agent in real-world, dynamic market environments.
+5.3 Ablation Studies
+We conduct ablation studies in the Arena to evaluate the impact of input formats and the reflection mechanism on agent performance.
+To further assess the effect of sequence length across textual and visual modalities, we perform controlled backtesting on NASDAQ.
+Backtesting mitigates dynamic data fluctuations, allowing sequence length to be isolated as the primary variable.
+Model Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
+w/o Reflection
+LLaMa-3 (Dubey et al., 2024) 4.09 0.686 2.270 62.50 0.196
+DeepSeek (Liu et al., 2024a) 27.31 2.248 2.755 92.21 1.530
+Qwen-2.5 (Yang et al., 2024) 30.37 2.680 1.459 93.75 1.703
+Gemini-1.5 (Reid et al., 2024) 14.32 1.221 2.778 83.84 1.729
+19.04 1.604 1.165 90.91 1.376
+23.96 1.981 1.334 100.00 1.485
+GPT-4o (Hurst et al., 2024) 13.04 1.139 1.929 54.55 0.590
+17.07 1.454 1.504 72.73 0.967
+26.18 2.157 2.058 90.91 1.048
+w/ Reflection
+LLaMa-3 (Dubey et al., 2024) 10.25 1.667 2.403 66.67 0.694
+DeepSeek (Liu et al., 2024a) 30.62 2.470 1.512 100.00 1.633
+Qwen-2.5 (Yang et al., 2024) 38.91 3.324 1.073 100.00 3.099
+Gemini-1.5 (Reid et al., 2024) 29.45 2.519 3.012 93.51 2.210
+37.01 2.911 1.235 100.00 2.357
+41.33 3.195 0.159 100.00 20.113
+GPT-4o (Hurst et al., 2024) 33.65 2.671 2.171 98.70 2.142
+35.76 2.821 0.678 100.00 4.159
+47.69 3.610 0.533 100.00 6.777
+Table 5: Performance Comparison with and without Reflection Using Textual and Visual Inputs.
+Agents using visual data consistently outperform those using text alone, while combining both modalities yields the strongest results.
+Textual vs. Visual Inputs.
+We compare agents using textual-only inputs with those leveraging visualizations (scatter plots, line charts, bar graphs).
+Non-vision LLMs receive text only without reflection, while vision-capable LLMs are evaluated under text-only, visual-only, and combined-input settings.
+Corresponding prompts are provided in Appendix˜C, and details of LLM choices are in Appendix˜D.
+Results in Table˜5 show that visual inputs substantially improve returns, with combined inputs achieving the best overall performance.
+Impact of Reflection.
+Incorporating the reflection module further enhances performance across all LLMs.
+For GPT-4o, TR improves from 33.65% with text-only to 47.70% with visual input, a relative gain of 41.7%.
+Beyond TR, reflection also boosts SR and WR, indicating higher profitability, stability, and consistency in trading outcomes.
+These findings confirm that strategic reflection, particularly with visual data, strengthens agents’ reasoning and decision-making under volatile market conditions.
+Input Window TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
+Textual 5 days 3.63 0.102 1.588 62.5 0.064
+10 days 5.58 0.145 1.334 60.0 0.109
+15 days 6.47 0.171 1.681 57.5 0.102
+20 days 1.67 0.050 1.296 62.5 0.039
+Visual 5 days 8.21 0.207 1.396 57.5 0.149
+10 days 6.90 0.174 1.164 67.5 0.149
+15 days 6.90 0.174 1.216 60.0 0.143
+20 days 2.81 0.075 1.101 62.5 0.068
+Textual
+Visual
+5 days 14.33 0.372 1.719 60.0 0.216
+10 days 15.99 0.450 1.110 59.0 0.348
+15 days 14.73 0.380 1.684 57.5 0.226
+20 days 13.51 0.352 1.577 65.0 0.223
+Table 6: Performance with Textual and Visual Inputs Across Different Window Sizes.
+Window denotes the number of consecutive trading days provided as historical input before the current trading day.
+Textual + Visual consistently achieves the best performance, particularly with a 10-day window, highlighting the importance of both modality choice and time-window design.
+Impact of Time-Series Length.
+Experiments with GPT-4o as the backbone are summarized in Table˜6, which reports performance across modalities and time windows.
+Overall, “Textual + Visual” consistently outperforms single modalities, particularly with a 10-day window, achieving the highest TR (15.99%), SR (0.348), and Mean (0.450%).
+Visual inputs perform best in the 5–10 day range, offering strong returns, stability, and win rates.
+Performance declines at 20 days, likely due to noise accumulation or overfitting.
+These results underscore the importance of both multi-modal fusion and careful time-window selection for robust time-series decision-making.
 6 Conclusion
-In this paper, we introduced the Agent Trading Arena, a zero-sum game designed to simulate complex economic systems and evaluate LLMs on numerical reasoning tasks.
-Our results show that while LLMs struggle with plain-text numerical data (algebraic reasoning), their performance significantly improves when presented with visual data (geometric reasoning).
-This highlights the advantage of visual representations in supporting numerical reasoning, particularly in complex scenarios.
-The integration of a reflection module further enhances model performance, allowing LLMs to analyze better and interpret data.
-We validated these findings on NASDAQ Stock dataset, demonstrating that LLMs excel in visual geometric reasoning tasks, suggesting that LLMs may perform better with visual numerical data than with textual numerical data.
-Overall, our work offers insights into the strengths and limitations of LLMs in dynamic numerical reasoning tasks, particularly in the context of geometry vs.
-algebra, and sets the foundation for future research on improving their performance in real-world, interdisciplinary challenges.
+We presented the Agent Trading Arena, a closed-loop, zero-sum simulation framework designed to model complex market dynamics and evaluate LLMs on numerical reasoning tasks.
+Our experiments show that while LLMs struggle with plain-text numerical data, they achieve significantly higher returns when supported by visual representations, underscoring the importance of charts and graphs in complex decision-making.
+Integrating a reflection module further improves performance by enabling iterative analysis and strategic refinement.
+Validation on both NASDAQ and CSI confirms that vision-enhanced LLM agents consistently outperform text-only counterparts without additional training.
+Overall, this work highlights both the capabilities and limitations of LLMs in numerical reasoning and establishes the Agent Trading Arena as a robust testbed for advancing research in adaptive, real-world financial applications.
 Limitations
-This study evaluates LLMs within a virtual stock trading environment, focusing on their performance in visual geometric reasoning.
-While this controlled setting limits the generalizability to other domains, it provides valuable insights into LLMs’ capabilities.
-The reliance on high-quality visualizations, reflection modules, and substantial computational resources may restrict applicability in resource-constrained environments.
-Future research can address these limitations by broadening the scope to include diverse reasoning tasks, optimizing computational requirements, and exploring alternative modalities for more robust and generalizable assessments.
+This study investigates the visual reasoning abilities of LLMs in a virtual stock trading environment.
+While such a controlled setup enables systematic analysis, it inevitably simplifies many aspects of real-world financial systems, including uncertainty, feedback loops, and human interactions.
+Bridging the gap between virtual environments and real-world applications therefore remains a significant challenge.
+A promising direction for future work is to explore how agents can be gradually integrated into real financial workflows, interacting with live data, adapting to real-time signals, and contributing meaningful decisions or strategies.
+This would move beyond passive evaluation toward active participation, where agents not only respond to environments but also generate novel behavioral data and insights to inform downstream analysis, model training, and system design.
+Expanding the task scope and incorporating more diverse modalities may further support the development of generalizable and robust evaluation frameworks.
 Acknowledgments
-This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University under Grant 2024BDIAA01.
+This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University, under Grant 2024BDIAA01.
 Appendix A Agent Trading Arena
-A.1 Agent Trading Arena Details
-As illustrated in Figure 5, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection.
+A.1 Overview
+As illustrated in Fig.˜7, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection.
 In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
@@
 Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks.
-The action generation process is outlined in Equation 1, with input prompts shown in Figure 6 and the corresponding outputs displayed in the adjacent figure.
+The action generation process is outlined in Eq.˜1, with input prompts shown in Fig.˜8 and the corresponding outputs displayed in the adjacent figure.
 $\begin{cases}A_{\mathrm{date}+1}^{t}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{if~}t\text{~is Iters},\\ A_{\mathrm{date}}^{t+1}=\Psi\left(\mathrm{Ins},Z_{\mathrm{date}}^{t},S_{\mathrm{date}}\right),&\text{otherwise},\end{cases}$ (1)
@@
 To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment.
-The function $\phi$ facilitates environmental interactions, as shown in Algorithm 1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment.
+The function $\phi$ facilitates environmental interactions, as shown in Algorithm˜1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment.
 Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares.
@@
     $\text{Cash}\leftarrow\text{Price}_{\text{curr}}\cdot Q$;
-             if $<P.\text{Cash}$ Cash then
+    if $\text{Cash}<P.\text{Cash}$ then
        $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
@@
     $\text{Hold}\leftarrow P(N)$;
-[…]
+    if $\text{Hold}\neq\text{None}$ then
        $Q_{N}\leftarrow\text{Hold}[\text{``Q''}]-Q$;
@@
           $Z\leftarrow\text{SubmitOrder}(O,N,t,\text{Price}_{\text{curr}},Q)$;
-[…]
+$\text{Market}(\text{date},\text{Persons})$;
 return Z;
@@
 The reflection model serves as long-term memory, enabling self-reflection and the consolidation of knowledge.
-Figure 7: Visualization of Stock Inputs and Corresponding Trading Strategy Outputs.
 A.1.4 Reflection
@@
 where $Z^{\prime}_{\mathrm{date}}$ represents the long-term memory, and $\mathrm{DAYS}$ denotes the maximum number of days.
-Together, short-term and long-term memory provide essential context for the agents.
+Together, short-term and long-term memory provides essential context for the agents.
 Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors.
 Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
-Figure 8: Overview of Complete Workflow.
-Appendix B NASDAQ Stock
+Appendix B NASDAQ and CSI
 This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA.
 These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes.
-The NASDAQ stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays.
+The NASDAQ dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays.
 This dataset reflects current market trends and serves as a timely foundation for our research.
 It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
-The training and testing periods for MACD, StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and our system are shown in Table 7.
-StockFormer and TimesNet require longer training periods, with their training datasets spanning a broader time range compared to the other models.
-In contrast, our system does not require training and relies solely on historical stock data for making trading decisions.
-Strategy Training Testing
-Start End Start End
-MACD 29/7/2024 30/8/2024 3/9/2024 29/10/2024
-StockFormer 3/7/2023 30/8/2024 3/9/2024 29/10/2024
-TimesNet 3/7/2023 30/8/2024 3/9/2024 29/10/2024
-Ours 16/8/2024 30/8/2024 3/9/2024 29/10/2024
-Table 7: Training and Testing Periods for Various Models on the NASDAQ Stock Dataset.
+This study also selects seven representative stocks from the Chinese stock market (CSI), covering various sectors with industry-leading companies: Kweichow Moutai (600519.SH), China Merchants Bank (600036.SH), CATL (300750.SZ), BYD (002594.SZ), LONGi Green Energy (601012.SH), ZTE (000063.SZ) and Mindray (300760.SZ).
+These stocks span consumer goods, finance, new energy, telecommunications, and healthcare, representing leading companies with strong market influence, stable performance, and innovative potential.
+For CSI, we conducted a one-year simulation from January 2 to December 31, 2024, using the same initial capital setup, while excluding weekends and official trading holidays.
+The data set is sourced from Baostock and includes daily records such as the opening price, closing price, highest price, lowest price, trading volume, and multiple technical indicators.
+This comprehensive data set provides a solid foundation for analyzing market trends and investment strategies within the CSI framework.
+The training and testing periods for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), StockMixer (Fan and Shen, 2024), and our system is shown in Table˜7.
+StockFormer, TimesNet, and StockMixer require longer training periods, and their training data sets span a broader time range compared to the other models.
+In contrast, our system does not require training and relies solely on historical stock data to make trading decisions.
+Strategy NASDAQ CSI
+Training Testing Training Testing
+Start End Start End Start End Start End
+Buy & Hold w/o 25/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 21/12/2023 29/12/2023 2/1/2024 31/12/2024
+SMA w/o 1/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
+ZMR w/o 10/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 8/12/2023 29/12/2023 2/1/2024 31/12/2024
+MACD w/o 4/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 4/12/2023 29/12/2023 2/1/2024 31/12/2024
+StockFormer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
+TimesNet w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
+StockMixer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
+Ours w/o 16/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
+Table 7: Training and Testing Periods for Various Models on NASDAQ and CSI.
 The testing period is set from September 3, 2024, to October 29, 2024, to prevent potential data leaks that could provide prior knowledge to the GPT-4o system.
 This timeframe ensures the fairness of the evaluation by mitigating biases from such leaks, thereby enhancing the reliability of the experimental outcomes.
-Appendix C Visualization Input
-Figure 7 illustrates the system’s input prompts and corresponding outputs during the strategy update process.
+We also select two stocks from the CSI dataset, each representing a different growth pattern, for simulation.
+As shown in Table˜8, our method consistently outperforms the average trend of the selected stocks.
+Trend LLMs TR $\uparrow$ SR $\uparrow$
+Average Min Max
+-5.86 -13.77 -0.21 GPT4o -0.50 -0.021
+Qwen 2.01 0.207
+LlaMa3 1.44 0.122
+11.37 -4.25 26.05 GPT4o 20.61 0.465
+Qwen 16.46 0.420
+LlaMa3 4.14 0.167
+Table 8: Performance on CSI Stocks with Different Growth Patterns. Our method consistently outperforms the average trend of both selected stocks.
+Appendix C Prompt Design Across Modalities
+The prompts used for stock analysis with text-only, vision-only, and combined inputs are shown below. Additional prompts for others can be found at: https://github.com/wekjsdvnm/Agent-Trading-Arena.
+Text-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge.
+- You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation.
+- Your overall objective is to make as much profit as possible.
+Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B.
+This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B.
+Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns.
+Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit.
+Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information.
+The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy.
+In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) Output the response to the prompt above in JSON.
+Each analysis result should be started with "-", and ended with a line break.
+Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
+Visual-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge.
+- You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation.
+- Your overall objective is to make as much profit as possible.
+Stock information: [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B.
+This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B.
+Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns.
+Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver consistent profits with minimal risk.
+As traders weigh their options, strategic portfolio rebalancing tips towards these dividends-rich stocks seem to be the talk of the town.
+However, with the volatile nature of stock rumors, it remains essential to tread carefully and keep an eye on reliable sources for confirmation.
+Will this buzz hold substance, or is it merely hot air?
+Only time will reveal the truth behind the whispers, leaving the market in suspense.
+Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - C: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT Investment strategy: try to maximize profit.
+Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information.
+The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy.
+In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations.
+Output the response to the prompt above in JSON.
+Each analysis result should be started with "-", and ended with a line break.
+Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
+Combined-Input Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge.
+- You will be given information related to specific stocks, the whole market, and your existing…… Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B.
+This merger is rumored to lead to significant synergies…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit.
+Task: - Based on (but not limited to) the following information, …… describe the relationship between its content and investment strategy.
+In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations.
+Output the response to the prompt above in JSON.
+Each analysis result should be started with "-",…… Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
+Appendix D Simulation Process
+The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek (Liu et al., 2024a), Qwen-2.5 (Yang et al., 2024), Qwen-VL-32k (Wang et al., 2024b), Qwen-VL-128kk (Wang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, DeepSeek-chat, Qwen2.5-72B-Instruct, qwen-vl-max, qwen-vl-max-latest, and Gemini-1.5-pro.
+Appendix E Visualization Input
+Fig.˜9 illustrates the system’s input prompts and corresponding outputs during the strategy update process.
 The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
-Appendix D Simulation Process
+Figure 9: Visualization of Stock Inputs and Corresponding Trading Strategy Outputs.
+Appendix F Simulation Process
 In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis.
@@
 Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory.
-This entire process is illustrated in Figure 8.
-Appendix E Simulation Process
-The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), DeepSeek (Liu et al., 2024), Qwen-2.5 (Yang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, DeepSeek-chat, Qwen2.5-72B-Instruct, and Gemini-1.5-pro.
+This entire process is illustrated in Fig.˜10.
+Figure 10: Overview of Complete Workflow.
+Impact of Modality on LLM Competitiveness.
+We employed a relative evaluation method for this experiment.
+The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline.
+This setup aimed to explore the impact of different agents and modalities on LLM performance.
+The results are shown in Table˜9.
+The findings indicate that DeepSeek (Liu et al., 2024a) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
+LLMs T V TR $\uparrow$ SR $\uparrow$
+LLaMa-3 100.00 100.00
+DeepSeek + 35.03 + 45.63
+Gemini-1.5 + 10.99 + 14.93
++ 11.49 + 29.47
+GPT-4o + 10.39 + 28.29
++ 17.18 + 40.77
+Table 9: Performance Comparison in Trading Decisions w/o Reflection, Competing Pairwise with LLaMa-3.
+DeepSeek may possess unique strengths or optimizations that allow it to better adapt to the task’s complexities when competing with other models.